# Forecasting CPI inflation under economic policy and geo-political uncertainties

### v1

@@ 5.3 Experimental results and baseline comparison
Country Metrics DeepAR ARNNx NBeats ARFIMAx SARIMAx ARIMAx LR ARDL XGBoost TFT WARIMAx Proposed
FEWNet
Brazil RMSE 3.41 2.23 2.95 4.67 2.37 2.69 2.19 3.26 2.40 5.41 3.13 1.30
MASE 5.24 3.32 4.41 7.49 3.77 4.29 3.45 5.57 3.62 8.51 4.88 1.80
SMAPE(%) 37 25 34 68 29 33 26 52 27 85 39 12
Theil’s $U_{1}$ 0.20 0.14 0.22 0.38 0.16 0.19 0.14 0.25 0.17 0.48 0.23 0.09
MDRAE 4.42 2.45 5.22 8.38 3.84 4.44 3.70 5.24 4.41 9.26 5.45 1.97
MDAPE 0.30 0.20 0.34 0.59 0.27 0.30 0.23 0.40 0.26 0.46 0.39 0.10
Russia RMSE 2.00 4.16 2.49 12.68 6.65 7.80 4.45 3.65 6.21 3.87 3.72 0.88
MASE 4.66 10.17 5.63 33.84 17.62 20.26 11.28 9.67 16.06 9.96 9.44 2.21
SMAPE(%) 26 46 38 100 69 76 48 44 63 81 44 13
Theil’s $U_{1}$ 0.15 0.25 0.23 0.50 0.34 0.38 0.26 0.22 0.33 0.42 0.23 0.07
MDRAE 4.23 10.20 5.00 40.54 18.31 23.76 12.70 12.14 15.49 7.57 9.15 2.88
MDAPE 0.33 0.51 0.32 1.97 1.05 1.22 0.71 0.57 0.99 0.59 0.56 0.11
India RMSE 3.55 1.69 1.09 1.86 2.27 2.31 2.46 1.88 1.69 1.67 1.19 0.88
MASE 6.61 2.68 1.82 2.92 3.72 3.88 4.19 2.96 3.07 2.35 2.00 1.41
SMAPE(%) 52 33 22 28 51 54 58 38 38 24 23 15
Theil’s $U_{1}$ 0.27 0.21 0.12 0.17 0.29 0.30 0.33 0.23 0.21 0.16 0.13 0.09
MDRAE 9.18 2.18 1.70 2.43 3.91 5.16 4.26 2.37 3.93 1.86 2.46 1.22
MDAPE 0.68 0.26 0.20 0.38 0.30 0.34 0.44 0.30 0.36 0.15 0.22 0.15
China RMSE 1.33 2.01 0.90 10.45 4.54 5.21 2.35 3.34 0.99 2.11 4.85 1.70
MASE 1.73 2.76 1.18 15.68 6.68 7.67 3.24 4.75 1.25 2.77 7.18 1.93
SMAPE(%) 94 111 116 171 146 150 115 132 99 153 149 85
Theil’s $U_{1}$ 0.42 0.49 0.34 0.83 0.68 0.71 0.53 0.62 0.35 0.80 0.70 0.43
MDRAE 1.29 2.60 1.28 15.94 7.27 8.04 3.41 5.17 1.25 3.95 7.38 1.75
MDAPE 1.66 1.70 0.87 14.68 6.80 7.67 2.26 5.07 0.93 1.99 7.21 0.99
@@
Country Metrics DeepAR ARNNx NBeats ARFIMAx SARIMAx ARIMAx LR ARDL XGBoost TFT WARIMAx Proposed
FEWNet
Brazil RMSE 6.16 3.56 3.11 4.27 4.13 4.49 3.67 2.86 2.56 3.58 4.11 2.31
MASE 11.45 6.68 5.10 6.04 7.67 8.27 6.93 5.23 3.62 6.11 5.93 4.12
SMAPE(%) 91 57 47 58 64 69 58 47 40 55 57 39
Theil’s $U_{1}$ 0.45 0.27 0.29 0.47 0.32 0.36 0.27 0.23 0.21 0.31 0.45 0.19
MDRAE 12.66 6.81 5.60 3.67 9.31 8.31 7.49 5.87 5.11 6.85 3.76 4.84
MDAPE 0.86 0.60 0.48 0.36 0.56 0.65 0.70 0.44 0.41 0.50 0.39 0.36
Russia RMSE 3.26 6.12 5.26 11.69 9.90 9.49 6.97 6.67 2.16 2.72 5.22 2.24
MASE 8.39 18.32 14.63 35.12 28.78 27.25 20.31 19.75 5.10 7.00 15.03 5.28
SMAPE(%) 51 81 66 112 101 98 84 84 36 44 72 27
Theil’s $U_{1}$ 0.28 0.38 0.34 0.54 0.51 0.50 0.42 0.40 0.25 0.24 0.35 0.18
MDRAE 8.29 21.94 15.12 41.21 30.43 26.85 24.53 24.47 4.76 7.21 17.53 3.51
MDAPE 0.44 1.33 1.09 2.49 2.02 1.87 1.38 1.47 0.35 0.43 1.14 0.38
India RMSE 1.86 2.22 4.18 2.72 2.16 2.02 2.36 2.53 1.95 6.17 2.35 1.04
MASE 2.68 2.90 5.83 4.20 3.02 2.83 3.61 2.98 2.80 9.48 3.39 1.49
SMAPE(%) 26 35 46 38 34 31 47 37 33 65 37 17
Theil’s $U_{1}$ 0.15 0.23 0.29 0.21 0.21 0.19 0.26 0.26 0.20 0.37 0.21 0.10
MDRAE 3.25 4.31 6.22 6.46 4.02 4.07 4.21 2.12 4.28 13.47 3.70 1.42
MDAPE 0.31 0.24 0.49 0.42 0.30 0.28 0.41 0.21 0.26 1.07 0.40 0.14
China RMSE 1.71 1.85 1.43 18.37 3.48 4.04 1.59 2.00 2.87 2.05 7.40 2.01
MASE 1.95 2.15 1.75 25.72 4.23 5.05 1.89 2.33 3.12 2.42 10.17 2.37
SMAPE(%) 81 77 77 168 102 109 75 79 115 100 139 82
Theil’s $U_{1}$ 0.40 0.34 0.27 0.81 0.48 0.51 0.30 0.34 0.65 0.46 0.64 0.36
MDRAE 1.53 2.32 1.60 26.02 4.43 4.99 2.09 2.51 2.68 2.32 10.88 2.50
MDAPE 0.57 0.85 0.63 11.56 2.05 2.41 0.55 0.99 1.14 0.84 4.84 0.84
@@ 5.4 Statistical significance of the results
In this section, we assess the effectiveness of different forecasting models in terms of the statistical significance of measurement error differences using the model-agnostic multiple comparisons with the best (MCB) procedure [65]. This non-parametric test computes the average rank for each of the $\mathcal{M}$ forecasters based on their ex-ante accuracy across $\mathcal{D}$ datasets and identifies the model with the minimum average rank as the ‘best’ performing technique. The MCB test determines critical distances (CD) for each algorithm as $\Theta_{\alpha}\sqrt{\frac{\mathcal{M}\left(\mathcal{M}+1\right)}{6\mathcal{D}}}$, where $\Theta_{\alpha}$ is the critical value of the Tukey distribution at level $\alpha$, and it considers the CD of the best-performing model as the reference value for the test. Following the MCB test, we calculate the average rank of the algorithm and present the results of the MCB test for the RMSE metric in Figure 4. The plot reveals that the FEWNet framework has achieved the lowest mean rank of 2.12 among all competing algorithms for both forecast horizons. Therefore, it is deemed to be the best-performing model, followed by NBeats (4.38), XGBoost (4.50), ARNNx (5.25), and so forth. Furthermore, the upper limit of the critical distance for the FEWNet model (shown by the shaded area in the graphs) acts as the reference value for the test. When comparing SARIMAx, ARIMAx, and ARFIMAx to the baseline, it is observed that their critical intervals are much higher than the reference value and do not coincide with FEWNet. This indicates that their forecast performance is notably worse than the FEWNet approach. Our analysis underscores the statistical significance of the performance difference and the superiority of the proposed FEWNet model across datasets and forecast horizons. Moreover, the test establishes that FEWNet consistently and significantly outperforms other baseline models, as well as various machine learning or deep learning models.
@@
Figure 4: Visualization of the multiple comparisons with best (MCB) analysis for (a) Brazil, (b) Russia, (c) India (d) China [BRIC] countries: In the figure, for example in (a) FEWNet- 2.12 means that the average rank of the proposed algorithm FEWNet based on RMSE error metric is 2.12, the same explanation applies to other countries and algorithms

### v2

@@ 5.3 Experimental results and baseline comparison
Country Metrics RW RWD AR MSGARCH DeepAR ARNNx NBeats ARFIMAx SARIMAx ARIMAx LR ARDL XGBoost TFT WARIMAx Proposed FEWNet Brazil RMSE 4.13 4.29 3.80 3.59 3.41 2.23 2.95 4.67 2.37 2.69 2.19 3.26 2.40 5.41 3.13 1.30 MASE 6.16 6.40 5.58 5.19 5.24 3.32 4.41 7.49 3.77 4.29 3.45 5.57 3.62 8.51 4.88 1.80 SMAPE(%) 52 55 45 41 37 25 34 68 29 33 26 52 27 85 39 12 Theil’s $U_{1}$ 0.33 0.35 0.30 0.28 0.20 0.14 0.22 0.38 0.16 0.19 0.14 0.25 0.17 0.48 0.23 0.09 MDRAE 7.18 7.44 6.59 5.36 4.42 2.45 5.22 8.38 3.84 4.44 3.70 5.24 4.41 9.26 5.45 1.97 MDAPE 0.47 0.49 0.43 0.36 0.30 0.20 0.34 0.59 0.27 0.30 0.23 0.40 0.26 0.46 0.39 0.10 Russia RMSE 2.14 2.49 0.99 1.14 2.00 4.16 2.49 12.68 6.65 7.80 4.45 3.65 6.21 3.87 3.72 0.88 MASE 5.04 5.87 1.95 2.31 4.66 10.17 5.63 33.84 17.62 20.26 11.28 9.67 16.06 9.96 9.44 2.21 SMAPE(%) 33 40 11 13 26 46 38 100 69 76 48 44 63 81 44 13 Theil’s $U_{1}$ 0.19 0.23 0.08 0.09 0.15 0.25 0.23 0.50 0.34 0.38 0.26 0.22 0.33 0.42 0.23 0.07 MDRAE 4.00 4.83 1.83 2.14 4.23 10.20 5.00 40.54 18.31 23.76 12.70 12.14 15.49 7.57 9.15 2.88 MDAPE 0.27 0.33 0.08 0.12 0.33 0.51 0.32 1.97 1.05 1.22 0.71 0.57 0.99 0.59 0.56 0.11 India RMSE 0.99 0.91 1.23 2.18 3.55 1.69 1.09 1.86 2.27 2.31 2.46 1.88 1.69 1.67 1.19 0.88 MASE 1.53 1.34 2.09 3.68 6.61 2.68 1.82 2.92 3.72 3.88 4.19 2.96 3.07 2.35 2.00 1.41 SMAPE(%) 16 15 21 33 52 33 22 28 51 54 58 38 38 24 23 15 Theil’s $U_{1}$ 0.10 0.09 0.12 0.19 0.27 0.21 0.12 0.17 0.29 0.30 0.33 0.23 0.21 0.16 0.13 0.09 MDRAE 1.51 0.81 2.66 4.90 9.18 2.18 1.70 2.43 3.91 5.16 4.26 2.37 3.93 1.86 2.46 1.22 MDAPE 0.15 0.11 0.23 0.35 0.68 0.26 0.20 0.38 0.30 0.34 0.44 0.30 0.36 0.15 0.22 0.15 China RMSE 1.58 1.71 0.82 1.75 1.33 2.01 0.90 10.45 4.54 5.21 2.35 3.34 0.99 2.11 4.85 1.70 MASE 2.02 2.19 1.07 2.40 1.73 2.76 1.18 15.68 6.68 7.67 3.24 4.75 1.25 2.77 7.18 1.93 SMAPE(%) 173 174 103 107 94 111 116 171 146 150 115 132 99 153 149 85 Theil’s $U_{1}$ 0.94 0.95 0.42 0.46 0.42 0.49 0.34 0.83 0.68 0.71 0.53 0.62 0.35 0.80 0.70 0.43 MDRAE 2.28 2.44 1.10 2.41 1.29 2.60 1.28 15.94 7.27 8.04 3.41 5.17 1.25 3.95 7.38 1.75 MDAPE 1.34 1.46 0.76 2.42 1.66 1.70 0.87 14.68 6.80 7.67 2.26 5.07 0.93 1.99 7.21 0.99
@@
Country Metrics RW RWD AR MSGARCH DeepAR ARNNx NBeats ARFIMAx SARIMAx ARIMAx LR ARDL XGBoost TFT WARIMAx Proposed FEWNet Brazil RMSE 3.56 3.93 2.73 2.78 6.16 3.56 3.11 4.27 4.13 4.49 3.67 2.86 2.56 3.58 4.11 2.31 MASE 5.34 5.89 4.34 4.64 11.45 6.68 5.10 6.04 7.67 8.27 6.93 5.23 3.62 6.11 5.93 4.12 SMAPE(%) 50 57 40 42 91 57 47 58 64 69 58 47 40 55 57 39 Theil’s $U_{1}$ 0.37 0.43 0.25 0.25 0.45 0.27 0.29 0.47 0.32 0.36 0.27 0.23 0.21 0.31 0.45 0.19 MDRAE 5.11 4.52 2.63 2.97 12.66 6.81 5.60 3.67 9.31 8.31 7.49 5.87 5.11 6.85 3.76 4.84 MDAPE 0.41 0.40 0.38 0.42 0.86 0.60 0.48 0.36 0.56 0.65 0.70 0.44 0.41 0.50 0.39 0.36 Russia RMSE 2.06 2.74 1.03 1.82 3.26 6.12 5.26 11.69 9.90 9.49 6.97 6.67 2.16 2.72 5.22 2.24 MASE 4.87 6.40 2.67 4.63 8.39 18.32 14.63 35.12 28.78 27.25 20.31 19.75 5.10 7.00 15.03 5.28 SMAPE(%) 34 48 20 32 51 81 66 112 101 98 84 84 36 44 72 27 Theil’s $U_{1}$ 0.23 0.33 0.10 0.16 0.28 0.38 0.34 0.54 0.51 0.50 0.42 0.40 0.25 0.24 0.35 0.18 MDRAE 4.58 5.80 2.12 4.71 8.29 21.94 15.12 41.21 30.43 26.85 24.53 24.47 4.76 7.21 17.53 3.51 MDAPE 0.33 0.43 0.20 0.27 0.44 1.33 1.09 2.49 2.02 1.87 1.38 1.47 0.35 0.43 1.14 0.38 India RMSE 3.72 3.73 2.92 1.92 1.86 2.22 4.18 2.72 2.16 2.02 2.36 2.53 1.95 6.17 2.35 1.04 MASE 6.20 6.21 4.87 2.77 2.68 2.90 5.83 4.20 3.02 2.83 3.61 2.98 2.80 9.48 3.39 1.49 SMAPE(%) 51 51 43 26 26 35 46 38 34 31 47 37 33 65 37 17 Theil’s $U_{1}$ 0.26 0.26 0.21 0.15 0.15 0.23 0.29 0.21 0.21 0.19 0.26 0.26 0.20 0.37 0.21 0.10 MDRAE 10.76 10.80 8.25 3.71 3.25 4.31 6.22 6.46 4.02 4.07 4.21 2.12 4.28 13.47 3.70 1.42 MDAPE 0.70 0.70 0.54 0.24 0.31 0.24 0.49 0.42 0.30 0.28 0.41 0.21 0.26 1.07 0.40 0.14 China RMSE 3.15 3.23 2.23 2.08 1.71 1.85 1.43 18.37 3.48 4.04 1.59 2.00 2.87 2.05 7.40 2.01 MASE 3.95 4.06 2.74 2.63 1.95 2.15 1.75 25.72 4.23 5.05 1.89 2.33 3.12 2.42 10.17 2.37 SMAPE(%) 101 102 89 89 81 77 77 168 102 109 75 79 115 100 139 82 Theil’s $U_{1}$ 0.45 0.46 0.36 0.35 0.40 0.34 0.27 0.81 0.48 0.51 0.30 0.34 0.65 0.46 0.64 0.36 MDRAE 3.87 4.07 2.61 2.42 1.53 2.32 1.60 26.02 4.43 4.99 2.09 2.51 2.68 2.32 10.88 2.50 MDAPE 1.68 1.73 1.14 0.95 0.57 0.85 0.63 11.56 2.05 2.41 0.55 0.99 1.14 0.84 4.84 0.84
@@ 5.4 Statistical significance of the results
In this section, we assess the effectiveness of different forecasting models in terms of the statistical significance of measurement error differences using the model-agnostic multiple comparisons with the best (MCB) procedure Koning et al. (2005). This non-parametric test computes the average rank for each of the $\mathcal{M}$ forecasters based on their ex-ante accuracy across $\mathcal{D}$ datasets and identifies the model with the minimum average rank as the ‘best’ performing technique. The MCB test determines critical distances (CD) for each algorithm as $\Theta_{\alpha}\sqrt{\frac{\mathcal{M}\left(\mathcal{M}+1\right)}{6\mathcal{D}}}$, where $\Theta_{\alpha}$ is the critical value of the Tukey distribution at level $\alpha$, and it considers the CD of the best-performing model as the reference value for the test. Following the MCB test, we calculate the algorithm’s average rank and present the MCB test results for the RMSE metric in Figure 4. The plot reveals that the FEWNet framework has achieved the lowest mean rank of 2.75 among all competing algorithms for both forecast horizons. Therefore, it is deemed to be the best-performing model, followed by AR (5.75), XGBoost (6.25), MSGARCH (6.38), NBeats (6.50), and so forth. Furthermore, the upper limit of the critical distance for the FEWNet model (shown by the shaded area in the graphs) acts as the reference value for the test. When comparing SARIMAx, ARIMAx, and ARFIMAx to the baseline, it is observed that their critical intervals are much higher than the reference value and do not coincide with FEWNet. This indicates that their forecast performance is notably worse than the FEWNet approach. AR model obtained second place and fell short of FEWNet’s performance. MSGARCH, although designed for handling volatility, does not surpass FEWNet. Our analysis underscores the statistical significance of the performance difference and the superiority of the proposed FEWNet model across datasets and forecast horizons. Moreover, the test establishes that FEWNet consistently and significantly outperforms other baseline models, as well as various machine learning or deep learning models.
@@
Figure 4: Visualization of the multiple comparisons with best (MCB) analysis for BRIC countries. In the figure, for example, FEWNet - 2.75 means that the average rank of the proposed algorithm FEWNet based on the RMSE error metric is 2.75; the same explanation applies to other algorithms.
In addition to the MCB test, the Giacomini and Rossi (GR) test Giacomini & Rossi (2010) is employed to assess the robustness and forecasting ability of the proposed FEWNet model under “unstable environments” in comparison to other baseline models. The GR test is designed to evaluate whether one forecasting model consistently outperforms another over a specified period, particularly in data with structural breaks or regime shifts, which traditional tests like the Diebold-Mariano test might not effectively handle Diebold & Mariano (2002); Rossi & Sekhposyan (2016). GR test specifically examines the stability of relative local forecasting performance over time. We used the ‘murphydiagram’ package in R to perform the GR test. Based on the MCB test results, the top two baseline models identified are AR and XGBoost. We employ the GR test to further evaluate the stability of the relative predictive accuracy of these models over time. This test facilitates pairwise comparisons between FEWNet and each of the two baseline models, enabling a rigorous assessment of their relative performance. GR test is performed at a 10% level of significance ($\alpha$ = 0.90), with critical value (CV) thresholds indicated by thick black line in the plots. The fluctuation test diagrams display varying widths of CV bounds and different endpoints for the rolling window across time. This variability can be attributed to the choice of the parameter ($\mu$), which captures the size of the rolling window relative to the evaluation sample during GR test implementation. The null hypothesis of the fluctuation test (GR test) posits that both forecasting methods perform equally well (same expected score) at all time points, while the alternative hypothesis seeks to identify if their performance differs at any time point. The objective is to find the value of $\mu$ that helps pinpoint where the performance divergence occurs. Consequently, different forecasting methods result in different CV thresholds and rolling window sizes, reflecting the unique points of performance variation over time. These charts are based on long-term (24 months) forecasting horizons (refer to Figure 5). From the GR test, we can conclude that the proposed FEWNet model consistently outperforms the top-performing baseline models (AR and XGBoost) across the BRIC countries during test periods. These findings collectively establish FEWNet as a more accurate and stable forecasting model under dynamic economic conditions, highlighting its potential as a valuable tool for economic forecasting and policy analysis.
Figure 5: Results of the GR test (Giacomini & Rossi Test) of forecasting ability of the FEWNet with baselines (AR (left) and XGBoost (right)) with ($\alpha$=0.90 and $\mu$=1) for (a) Brazil, (b) Russia, (c) India, and (d) China. Points outside the CV have been marked under the red dotted circles representing that FEWNet and other two top performing benchmarks differ significantly.
5.5 Empirical results on ERM
We investigate the empirical risk of the proposed FEWNet model on BRIC datasets (training series). Empirically, ERM helps to improve the model’s accuracy by minimizing the empirical risk and reducing the level of error or deviation between the predicted and actual values of the dataset. The key challenge in the inflation series is to learn fast dynamics and to cancel the noise simultaneously. We tackle this problem by iterative application of wavelet decomposition using low and band-pass filters, resulting in a trend series and a hierarchy of detail series that contains the system’s dynamics at different scales. Several ARNNx are trained to model these decomposed series and provide forecasts of the original time series by combining individual forecasts. We theoretically have proved in Section 4 that using the FEWNet method (incorporated wavelet decomposition) reduces the empirical risk under some conditions. We now carry out practical experiments on BRIC countries’ datasets to show that the use of the wavelet coefficients is robust to model estimation and reduces the generalization risk for the majority of the training samples as depicted in Table 7. It is clear from Table 7 (in-sample predictions) that FEWNet has improved performance for highly non-stationary time series as compared to standard neural networks and other benchmarks also for out-of-sample forecasts as depicted in Tables 5 and 6. This has a series of practical implications and is important in policymaking. In an ideal situation, FEWNet will always perform better than the autoregressive neural network model. Also, as FEWNet can separate the fast dynamics from the slow ones and generate a combined forecast, it escapes the problem of underfitting and overfitting. Thus, FEWNet is more suitable for countries with complex, non-stationary CPI inflation time series and will be very useful for policymakers at central banks in BRIC countries to easily visualize the future path of inflation. A more detailed discussion on policy implications is given in Section 6.
Table 7: Empirical risk comparison of two different training periods (best results are made bold).
Training Period Models Brazil Russia India China
2003-01 to 2020-11 ARNNx 2.61 16.62 0.51 1.16
FEWNet 1.14 0.77 1.38 1.10
2003-01 to 2019-11 ARNNx 5.79 17.66 0.54 1.23
FEWNet 4.60 13.13 0.93 1.01
5.6 Robustness and sensitivity analysis
This section provides a more comprehensive robustness and sensitivity analysis of the FEWNet method. For instance, we first study the impact of HP and CF filters in the FEWNet model. This is of particular interest since Hamilton (2018) criticized the HP filter for generating spurious dynamic relations and producing end-sample values that differ significantly from mid-sample values. Drehmann & Yetman (2018) rejected this criticism by highlighting scenarios in macroeconomic contexts where the HP filter is useful and robust. HP filter is more accurate when applied to medium and long-term cycles instead of short-term fluctuations. We employ an empirical approach to verify the importance of HP and CF economic filters in the FEWNet model. HP filter is used to extract trend components, whereas CF filter is employed to isolate cyclical components from the CPI inflation, EPU, and GPRC series. We examined the relevance of HP and CF filters by comparing the experimental results of the FEWNet model with and without economic filters. Tables 8 and 9 summarize the performance of FEWNet and FEWNet without HP and CF filters (referred to as EWNet) for semi-long-run and long-run. The results indicate that these economic filters play a critical role in CPI inflation series analysis. The tables underscore the substantial improvement in predictive accuracy and robustness of the FEWNet model when economic filters are applied. This finding highlights the importance of integrating trend and cycle components using HP and CF filters, which enhances the model’s ability to capture the underlying dynamics of economic series. Consequently, the FEWNet model with economic filters proves to be a more effective tool for long-term economic forecasting in the BRIC countries, addressing and overcoming the traditional criticisms of using the HP filter.
In addition, we analyze the sensitivity of the experimental results relative to the choice of different wavelet filters and the level of wavelet decomposition. The selection of wavelet filters is crucial in the overarching design of the proposed FEWNet architecture. The model employs the “haar” wavelet transform filter to decompose the underlying CPI inflation series into detailed and smooth components. In order to rationalize the choice of the wavelet transform filter, we conduct an iterative out-of-sample forecast evaluation, comparing the performance of these competing models (FEWNet Model with different wavelet filters) over the semi-long-term (12 months) and long-term (24 months) forecast horizon. A grid-search approach is used to determine the optimal set of hyperparameters for the underlying ARNNx model within the FEWNet framework. The selection of the wavelet decomposition level ($\mathcal{K}$) is determined by the formula $\mathcal{K}=\lfloor\log_{e}(\mathcal{N})\rfloor$, where $\mathcal{N}$ represents the size of the training observations Percival & Walden (2000). This calculation yields $\mathcal{K}=5$. The same decomposition level is applied uniformly to ensure a fair comparison among different wavelet filters. It is standard practice in financial forecasting exercises to utilize 5 to 6 levels of detail. For monthly data, the first detail ($\delta_{1}$) captures oscillations between 2-4 months, $\delta_{2}$ between 4-8 months, $\delta_{3}$ between 8-16 months, $\delta_{4}$ between 16-32 months, and $\delta_{5}$ between 32-64 months. This approach effectively captures business cycle frequency dynamics, as additional details would require longer time period information, which is often unavailable for most macroeconomic series Souropanis & Vivian (2023). For a comprehensive comparison, we have considered four other wavelet transform filters, such as (a) Daubechies with length 8 (d8): known for its orthogonality and compact support, the Daubechies wavelet is effective in capturing intricate structures in the data; (b) Least asymmetric with length 8 (la8): this wavelet balances asymmetry and orthogonality, providing a more symmetric filter compared to Daubechies; (c) Coiflet with length 6 (c6): the Coiflet wavelet emphasizes both orthogonality and symmetry, making it suitable for analysis requiring smooth approximation components; (d) Best localized with length 14 (bl14): this wavelet is designed for optimal localization in both time and frequency domains, enhancing the detection of localized features in the data. The decomposition levels are uniformly set to five for all wavelet transform filters, utilizing the MODWT-based MRA approach.
Figure 6 showcases that the FEWNet model with the ‘haar’ wavelet filter consistently achieves the lowest mean rank across the BRIC countries, indicating its best-in-class performance among the evaluated wavelet transform filters. The rationale for the ‘haar’ wavelet’s superior performance lies in its simplicity and capability to handle abrupt changes in data. Unlike more complex wavelets, the ‘haar’ wavelet’s structure allows it to efficiently manage and adapt to the volatility inherent in economic series, particularly in the presence of regime shifts or exogenous shocks. This efficiency is reflected in its lower forecast errors and greater stability in predictions.
Table 8: Performance comparison between the proposed FEWNet model and their variants for 12 months ahead forecasts with exogenous factors EPU and GPRC (best results are made bold).
Country Metrics FEWNet FEWNet - d8 FEWNet - la8 FEWNet - c6 FEWNet - bl14 EWNet
Brazil RMSE 1.30 2.76 3.10 1.36 1.93 3.63
MASE 1.80 4.16 4.79 1.77 2.70 5.94
SMAPE(%) 12 32 39 12 20 54
Theil’s $\mathrm{U}_{1}$ 0.09 0.19 0.22 0.09 0.13 0.28
MDRAE 1.97 3.81 7.22 0.92 3.50 5.29
MDAPE 0.10 0.27 0.36 0.11 0.18 0.44
Russia RMSE 0.88 2.47 0.85 0.57 5.83 2.49
MASE 2.21 5.90 2.16 1.29 15.48 4.88
SMAPE(%) 13 40 13 8 63 32
Theil’s $U_{1}$ 0.07 0.23 0.06 0.04 0.31 0.22
MDRAE 2.88 4.99 2.13 1.24 16.64 3.42
MDAPE 0.11 0.29 0.14 0.07 1.00 0.21
India RMSE 0.88 1.41 2.73 2.15 3.59 1.77
MASE 1.41 2.45 5.14 3.48 6.62 3.34
SMAPE(%) 15 28 44 31 53 31
Theil’s $U_{1}$ 0.09 0.15 0.22 0.19 0.28 0.16
MDRAE 1.22 3.28 6.18 2.49 7.79 4.61
MDAPE 0.15 0.24 0.57 0.34 0.74 0.36
China RMSE 1.70 4.06 1.85 2.39 4.63 0.86
MASE 1.93 5.87 2.06 3.47 6.83 1.12
SMAPE(%) 85 138 91 123 148 105
Theil’s $U_{1}$ 0.43 0.67 0.54 0.53 0.69 0.36
MDRAE 1.75 7.06 1.16 3.57 6.20 1.00
MDAPE 0.99 5.23 1.62 3.61 7.08 1.04
Table 9: Performance comparison between the proposed FEWNet model and their variants for 24 months ahead forecasts with exogenous factors EPU and GPRC (best results are made bold).
Country Metrics FEWNet FEWNet - d8 FEWNet - la8 FEWNet - c6 FEWNet - bl14 EWNet
Brazil RMSE 2.31 1.92 4.06 3.42 2.82 2.65
MASE 4.12 3.30 7.59 6.04 4.66 4.52
SMAPE(%) 39 34 64 52 43 41
Theil’s $\mathrm{U}_{1}$ 0.19 0.15 0.31 0.24 0.25 0.23
MDRAE 4.84 3.14 9.31 6.12 5.51 5.12
MDAPE 0.36 0.22 0.71 0.65 0.40 0.41
Russia RMSE 2.24 4.84 2.94 3.56 3.81 2.90
MASE 5.28 12.96 7.09 9.53 9.92 6.71
SMAPE(%) 27 116 40 77 51 41
Theil’s $\mathrm{U}_{1}$ 0.18 0.55 0.24 0.39 0.28 0.23
MDRAE 3.51 12.46 6.29 11.08 6.88 7.61
MDAPE 0.38 0.87 0.36 0.72 0.52 0.45
India RMSE 1.04 2.32 3.57 1.96 3.44 2.35
MASE 1.49 3.20 5.78 2.87 5.03 2.88
SMAPE(%) 17 44 48 26 43 26
Theil’s $\mathrm{U}_{1}$ 0.10 0.25 0.25 0.16 0.25 0.19
MDRAE 1.42 3.55 9.58 3.85 6.44 3.24
MDAPE 0.14 0.28 0.59 0.30 0.48 0.20
China RMSE 2.01 2.43 2.32 1.90 2.22 3.76
MASE 2.37 3.00 2.75 2.14 2.68 4.69
SMAPE(%) 82 97 85 98 87 106
Theil’s $\mathrm{U}_{1}$ 0.36 0.45 0.36 0.37 0.37 0.49
MDRAE 2.50 2.46 2.83 1.86 2.37 5.06
MDAPE 0.84 0.99 0.91 0.71 1.07 2.18
Figure 6: Visualization of the MCB analysis for BRIC countries for different FEWNet variants of the wavelet filters along with the proposed FEWNet algorithm. In the figure, for example, FEWNet - 1.62 means that the average rank of the FEWNet based on the RMSE error metric is 1.62; the same explanation applies to its variants based on different wavelet filters.